A piece of iron of specific heat capacity 0.04JKg-1K
-1 and mass 400 kg, is quickly
dropped into 30 kg of water at 10 oC contain in a calorimeter of 120 kg and specific
heat capacity of 0.1 JKg-1K
-1
. If the temperature of the mixture is 30 oC, calculate the
initial temperature of the hot iron (4200 JKg-1K
-1 = specific heat capacity of water)
1
Expert's answer
2020-07-08T13:15:05-0400
"m_ic_i(T_i-T)=(m_wc_w+m_cc_c)(T-T_w)"

"(400)(400)(T_i-30)\\=(30(4200)+120(100))(30-10)\\T_i=47.25\degree C"